# Hierarchical filtrations of torsion-free sheaves and birational geometry

### v1

@@ 7.2. Minimal models and coding–theoretic optimization
We adopt the notation of the proposition. Let
$f:Y\to X$
be the contraction of a set of (1)(-1)-curves $E_{1},\dots,E_{m}\subset Y$, and let $\mathcal{E}_{Y}$ be a vector bundle of rank $r$ on $Y$. Let $D_{Y}\subset Y(\mathbb{F}_{q})$ be the evaluation set, with $D_{E}=D_{Y}\cap\bigcup_{i}E_{i}$ the points on exceptional divisors, $D_{E}=N_{E}$, and $D=f(D_{Y}\setminus D_{E})$ the evaluation set on the minimal model $X$, $D=N_{Y}-N_{E}$.
Step 1:
By Proposition 5.5, the global sections of $\mathcal{E}_{Y}$ descend via $f_{*}$ to global sections of the bundle $\mathcal{E}=f_{*}\mathcal{E}_{Y}$ on $X$:
$H^{0}(X,\mathcal{E})\cong H^{0}(Y,\mathcal{E}_{Y}).$
Hence the code dimensions are identical:
$\dim C_{X}=\dim\operatorname{ev}_{D}(H^{0}(X,\mathcal{E}))=\dim H^{0}(X,\mathcal{E})=\dim H^{0}(Y,\mathcal{E}_{Y})=\dim C_{Y}=k.$
Step 2:
Let $d_{\min}(C_{Y})$ denote the minimal Hamming weight of a nonzero codeword of $C_{Y}$. Every section of $\mathcal{E}_{Y}$ vanishes along the exceptional divisors $E_{i}$ by the twisting or by the kernel of the elementary transform along $E_{i}$. Consequently, for all $s\in H^{0}(Y,\mathcal{E}_{Y})$, the coordinates corresponding to $D_{E}$ are identically zero.
The restriction of $s$ to $D^{\prime}:=D_{Y}\setminus D_{E}$ defines a codeword in $C_{X}$. Conversely, any nonzero codeword in $C_{X}$ lifts to a codeword in $C_{Y}$ with zeros in the $D_{E}$ coordinates. Hence the minimum weight satisfies
$d_{\min}(C_{X})=d_{\min}(C_{Y})=:d_{\min}.$
Step 3:
By definition, the normalized minimum distance is the ratio of absolute minimum distance to total number of coordinates:
$\delta_{Y}:=\frac{d_{\min}}{r\,N_{Y}},\qquad\delta_{X}:=\frac{d_{\min}}{r\,(N_{Y}-N_{E})}.$
Since $N_{E}\geq 0$, we have $N_{Y}-N_{E}\leq N_{Y}$, so
$\delta_{X}=\frac{d_{\min}}{r(N_{Y}-N_{E})}\geq\frac{d_{\min}}{rN_{Y}}=\delta_{Y}.$
Step 4:
The relative improvement in normalized distance is
$\frac{\delta_{X}}{\delta_{Y}}=\frac{d_{\min}/[r(N_{Y}-N_{E})]}{d_{\min}/(rN_{Y})}=\frac{N_{Y}}{N_{Y}-N_{E}}>1,$
as long as $N_{E}>0$, and it grows with the proportion of evaluation points lying on exceptional divisors.
All statements (i)–(iv) of the proposition follow directly from the descent of global sections via $f_{*}$, the structure of sections vanishing on exceptional divisors, and the definitions of absolute and normalized minimum distance. This completes the proof. ∎
@@
Let $X=\mathbb{P}^{2}_{\mathbb{F}_{5}}$ be the projective plane over $\mathbb{F}_{5}$, and let $f:Y\to X$ be the blow–up of $X$ at the $\mathbb{F}_{5}$–rational point $p=[1:0:0]$, with exceptional divisor $E\subset Y$.
@@
where $H=f^{*}(\text{line})$ is the pullback of a line in $\mathbb{P}^{2}$. Then $H^{0}(Y,\mathcal{E}_{Y})$ has dimension
$h^{0}(Y,\mathcal{E}_{Y})=h^{0}(\mathbb{P}^{2},\mathcal{O}(3))-1=10-1=9,$
because every cubic passing through $p$ defines a section, and subtracting $E$ forces vanishing at the blown-up point.
@@
with $\dim C_{Y}(\mathcal{E}_{Y},D_{Y})=\dim C_{Y}(\mathcal{E}_{Y},D^{\prime})=9$ and minimum distance
@@
$\delta_{Y}=\frac{d_{\min}}{N_{Y}}=\frac{4}{12}=\frac{1}{3}.$
@@
On the contracted surface $X=\mathbb{P}^{2}$, the direct image bundle $\mathcal{E}:=f_{*}\mathcal{E}_{Y}=\mathcal{O}_{X}(3)$ has
[…]
Let $D:=f(D^{\prime})\subset X(\mathbb{F}_{5})$, with #D=10\#D=10. The corresponding code is
$C_{X}(\mathcal{E},D)=\operatorname{ev}_{D}(H^{0}(X,\mathcal{E}))\subset\mathbb{F}_{5}^{10}.$
The minimum distance remains $d_{\min}(C_{X}(\mathcal{E},D))=4$, and the dimension is still $9$. However, the normalized minimum distance is
$\delta_{X}=\frac{d_{\min}}{N}=\frac{4}{10}=0.40>0.33=\delta_{Y}.$
Conclusion.
Contracting the exceptional divisor removes the degenerate coordinates introduced by evaluation points on $E$. While the absolute dimension and absolute minimum distance remain unchanged, the relative minimum distance (distance per code length) improves from $0.33$ to $0.40$.
@@ Conclusion.
Let $X=\mathbb{P}^{2}_{\mathbb{F}_{7}}$ and let $f:Y\to X$ be the blowup of $X$ at the $\mathbb{F}_{7}$-rational point $p=[1:0:0]$. Let $E\subset Y$ denote the exceptional (1)(-1)-curve, with $E(\mathbb{F}_{7})$ containing 8 rational points.
@@
The global sections of each summand are:
[…]
$H^{0}(Y,\mathcal{O}_{Y}(2H-E))$ $\cong H^{0}(X,\mathcal{O}_{X}(2))-1=6-1=5,$
$H^{0}(Y,\mathcal{O}_{Y}(H-E))$ $\cong H^{0}(X,\mathcal{O}_{X}(1))-1=3-1=2.$
Hence
$\dim_{\mathbb{F}_{7}}H^{0}(Y,\mathcal{E}_{Y})=9+5+2=16.$
@@
Every section of $\mathcal{E}_{Y}$ vanishes along $E$ (due to the E-E twist) therefore the coordinates corresponding to $D_{E}$ are all zeros:
@@
The absolute minimum distance (number of nonzero coordinates) for $C_{Y}(\mathcal{E}_{Y},D_{Y})$ is
$d_{\min}(C_{Y}(\mathcal{E}_{Y},D_{Y}))=d_{\min}(C_{Y}(\mathcal{E}_{Y},D^{\prime}))=4\quad(\text{computed on }D^{\prime}).$
The normalized (relative) minimum distance is
$\delta_{Y}=\frac{d_{\min}}{3\cdot 18}=\frac{4}{54}\approx 0.074.$
@@
On the contracted surface $X$, the direct image bundle
$\mathcal{E}:=f_{*}\mathcal{E}_{Y}\cong\mathcal{O}_{X}(3)\oplus\mathcal{O}_{X}(2)\oplus\mathcal{O}_{X}(1)$
by Proposition 5.5. Let
@@
The corresponding code $C_{X}(\mathcal{E},D)$ is
$C_{X}(\mathcal{E},D)=\operatorname{ev}_{D}(H^{0}(X,\mathcal{E}))\subset(\mathbb{F}_{7}^{3})^{13}.$
Dimension remains $16$, absolute minimum distance remains $d_{\min}(C_{X}(\mathcal{E},D))=4$ and normalized minimum distance improves:
$\delta_{X}=\frac{4}{3\cdot 13}=\frac{4}{39}\approx 0.103>\delta_{Y}\approx 0.074.$
Consequently, contracting the exceptional divisor $E$ removes the 15 degenerate coordinates ($5$ points $\times$ 33 ranks), while keeping the dimension and absolute minimum distance unchanged. As a result, the relative minimum distance increases from $\sim 0.074$ to $\sim 0.103$, demonstrating quantitatively the effect described in Proposition 7.3.
This shows that the MMP operation (contraction of (1)(-1)-curves) can strictly improve the coding-theoretic quality of AG codes when degenerate coordinates are introduced by exceptional divisors.

### v3

@@ 7.2. Minimal models and coding–theoretic optimization
Immediate from Proposition 7.1 since removing zero coordinates preserves $d_{\min}$ while reducing length. ∎
@@
Let $X=\mathbb{P}^{2}_{\mathbb{F}_{5}}$ be the projective plane over $\mathbb{F}_{5}$, and let $f:Y\to X$ be the blow-up of $X$ at the $\mathbb{F}_{5}$–rational point $p=[1:0:0]$, with exceptional divisor $E\subset Y$.
@@
where $H=f^{*}(\text{line})$ is the pullback of a line in $\mathbb{P}^{2}$.
Global sections.
From the exact sequence
$0\to\mathcal{O}_{Y}(3H-E)\to\mathcal{O}_{Y}(3H)\to\mathcal{O}_{E}(3H|_{E})\to 0,$
where $E\cong\mathbb{P}^{1}$ and $\mathcal{O}_{E}(3H|_{E})\cong\mathcal{O}_{\mathbb{P}^{1}}(3)$, we have
$0\to H^{0}(Y,\mathcal{O}_{Y}(3H-E))\to H^{0}(Y,\mathcal{O}_{Y}(3H))\to H^{0}(E,\mathcal{O}_{\mathbb{P}^{1}}(3)).$
Since $H^{0}(Y,\mathcal{O}_{Y}(3H))\cong H^{0}(X,\mathcal{O}_{X}(3))$ has dimension $10$ and $H^{0}(E,\mathcal{O}_{\mathbb{P}^{1}}(3))$ has dimension $4$, we obtain
$h^{0}(Y,\mathcal{E}_{Y})\geq 10-4=6.$
In fact, equality holds as the map $H^{0}(Y,\mathcal{O}_{Y}(3H))\to H^{0}(E,\mathcal{O}_{\mathbb{P}^{1}}(3))$ is surjective (cubics can take arbitrary values at $p$ to order $3$). Thus
$h^{0}(Y,\mathcal{E}_{Y})=6.$
@@
with $\dim C_{Y}(\mathcal{E}_{Y},D_{Y})=\dim C_{Y}(\mathcal{E}_{Y},D^{\prime})=6$.
For a concrete minimum distance, note that a cubic vanishing at $p$ can vanish at up to $7$ additional $\mathbb{F}_{5}$-points on a line (since a line has 66 $\mathbb{F}_{5}$-points, and the cubic restricted to the line is a degree $3$ polynomial). A careful choice of evaluation points yields, for instance,
@@
$\delta_{Y}=\frac{d_{\min}}{N_{Y}}=\frac{4}{12}=\frac{1}{3}\approx 0.333.$
@@
On the contracted surface $X=\mathbb{P}^{2}$, the direct image bundle
$\mathcal{E}:=f_{*}\mathcal{E}_{Y}=\mathcal{O}_{X}(3)$
has $\dim H^{0}(X,\mathcal{E})=10$, but only the $6$-dimensional subspace corresponding to cubics vanishing at $p$ yields nonzero codes when evaluated away from $p$. Let
$D:=f(D^{\prime})\subset X(\mathbb{F}_{5}),\quad\#D=10.$
The corresponding code using the same $6$-dimensional subspace is
$C_{X}(\mathcal{E},D)=\operatorname{ev}_{D}(H^{0}(Y,\mathcal{E}_{Y}))\subset\mathbb{F}_{5}^{10},$
where we identify $H^{0}(Y,\mathcal{E}_{Y})$ with the subspace of $H^{0}(X,\mathcal{O}_{X}(3))$ vanishing at $p$.
The minimum distance remains $d_{\min}(C_{X}(\mathcal{E},D))=4$, and the dimension is still $6$. The normalized minimum distance is
$\delta_{X}=\frac{d_{\min}}{10}=\frac{4}{10}=0.40>0.333=\delta_{Y}.$
@@
Let $X=\mathbb{P}^{2}_{\mathbb{F}_{7}}$ and let $f:Y\to X$ be the blowup of $X$ at the $\mathbb{F}_{7}$-rational point $p=[1:0:0]$. Let $E\subset Y$ denote the exceptional (1)(-1)-curve, with $E\cong\mathbb{P}^{1}$ having 88 $\mathbb{F}_{7}$-rational points. Denote by $\mathcal{I}_{p}\subset\mathcal{O}_{X}$ the ideal sheaf of the point $p$.
@@
For each summand, we compute dimensions using that $H^{0}(Y,\mathcal{O}_{Y}(dH-E))$ is isomorphic to the space of degree $d$ homogeneous polynomials vanishing at $p$:
$h^{0}(Y,\mathcal{O}_{Y}(H-E))$ $=h^{0}(X,\mathcal{I}_{p}\otimes\mathcal{O}_{X}(1))=\dim\{\text{linear forms through }p\}=2,$
$h^{0}(Y,\mathcal{O}_{Y}(2H-E))$ $=h^{0}(X,\mathcal{I}_{p}\otimes\mathcal{O}_{X}(2))=\dim\{\text{quadrics through }p\}=5,$
$h^{0}(Y,\mathcal{O}_{Y}(3H-E))$ $=h^{0}(X,\mathcal{I}_{p}\otimes\mathcal{O}_{X}(3))=\dim\{\text{cubics through }p\}=9.$
The dimensions follow from: $h^{0}(\mathbb{P}^{2},\mathcal{O}(d))=\binom{d+2}{2}$ and imposing one independent linear condition reduces dimension by 1. Hence
$h^{0}(Y,\mathcal{E}_{Y})=9+5+2=16.$
@@
We assume $D^{\prime}$ consists of points not on $E$ and in general position.
@@
Since every section of $\mathcal{E}_{Y}$ vanishes along $E$ (due to the E-E twist), the coordinates corresponding to $D_{E}$ are identically zero:
@@
To estimate the minimum distance, note that a nonzero section $s=(s_{3},s_{2},s_{1})\in H^{0}(Y,\mathcal{E}_{Y})$ has components $s_{d}\in H^{0}(Y,\mathcal{O}_{Y}(dH-E))$. Each $s_{d}$ vanishes at $p$ and can vanish at additional points. For generic $D^{\prime}$, a section can vanish at at most $(\deg(s_{3})+\deg(s_{2})+\deg(s_{1}))-1=(3+2+1)-1=5$ points of $D^{\prime}$ (by a parameter count). Thus we may have
$d_{\min}(C_{Y}(\mathcal{E}_{Y},D_{Y}))=d_{\min}(C_{Y}(\mathcal{E}_{Y},D^{\prime}))=13-5=8.$
The normalized minimum distance is
$\delta_{Y}=\frac{d_{\min}}{3\cdot 18}=\frac{8}{54}\approx 0.148.$
@@
On $X=\mathbb{P}^{2}$, consider the subspace
$V:=H^{0}(X,\mathcal{I}_{p}\otimes\mathcal{O}_{X}(3))\oplus H^{0}(X,\mathcal{I}_{p}\otimes\mathcal{O}_{X}(2))\oplus H^{0}(X,\mathcal{I}_{p}\otimes\mathcal{O}_{X}(1))$
of $H^{0}(X,\mathcal{O}_{X}(3)\oplus\mathcal{O}_{X}(2)\oplus\mathcal{O}_{X}(1))$, which corresponds to $H^{0}(Y,\mathcal{E}_{Y})$ under the natural isomorphism $H^{0}(Y,\mathcal{E}_{Y})\cong V$. Let
@@
The code using this subspace is
$C_{X}(V,D)=\operatorname{ev}_{D}(V)\subset(\mathbb{F}_{7}^{3})^{13}.$
The dimension remains $16$, and the minimum distance analysis is similar: a nonzero section $(s_{3},s_{2},s_{1})\in V$ vanishes at $p$ and can vanish at up to 5 additional points of $D$ (since $s_{3}$ is a cubic, $s_{2}$ a quadric, $s_{1}$ a linear form, all vanishing at $p$). Thus
$d_{\min}(C_{X}(V,D))=13-5=8.$
The normalized minimum distance improves to
$\delta_{X}=\frac{8}{3\cdot 13}=\frac{8}{39}\approx 0.205>0.148\approx\delta_{Y}.$
Remark 7.8.
For a contraction $f:Y\to X$ of (1)(-1)-curves $E_{1},\dots,E_{m}$ and a bundle $\mathcal{E}_{Y}$ whose sections vanish along $\bigcup_{i}E_{i}$, let $N_{E}$ be the number of evaluation points on the exceptional locus. Then
$\frac{\delta_{X}}{\delta_{Y}}=\frac{N_{Y}}{N_{Y}-N_{E}}>1,$
provided $N_{E}<N_{Y}$. This shows the normalized distance improvement is purely combinatorial, depending only on the proportion of “wasted” evaluation points.